OpenStudy (anonymous):

Angle x lies in the 3rd quadrant and tanx=7/24. a_ Determine an exact value for cos2x

OpenStudy (anonymous):

Tanx is opposite over hypotenuse so using Pythagorean theorem we can use 24^2 -7^2 = x^2 and x =22.956 so now use Cos(2x) double angle theorem to do Cos(x)^2 - sin(x) ^2 now that you have the whole triangle